# Radar-Guided Polynomial Fitting for Metric Depth Estimation

### v1

@@ Abstract
We propose PolyRad, a novel radar-guided depth estimation method that introduces polynomial fitting to transform scaleless depth predictions from pretrained monocular depth estimation (MDE) models into metric depth maps. Unlike existing approaches that rely on complex architectures or expensive sensors, our method is grounded in a simple yet fundamental insight: using polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust depth predictions non-uniformly across depth ranges. Although MDE models often infer reasonably accurate local depth structure within each object or local region, they may misalign these regions relative to one another, making a linear scale-and-shift transformation insufficient given three or more of these regions. In contrast, PolyRad generalizes beyond linear transformations and is able to correct such misalignments by introducing inflection points. Importantly, our polynomial fitting framework preserves structural consistency through a novel training objective that enforces monotonicity via first-derivative regularization. PolyRad achieves state-of-the-art performance on the nuScenes, ZJU-4DRadarCam, and View-of-Delft datasets, outperforming existing methods by 30.3% in MAE and 37.2% in RMSE.
@@ 1 Introduction
Metric 3D reconstruction is critical for spatial tasks such as self-driving [34, 14], where it is necessary for one to perceive the structure of the 3D environment in order to navigate. In many such systems, multiple sensors—including cameras, lidar, and radar provide complementary information about the 3D scene. While lidar sensors offer dense, accurate point clouds, they are expensive and not widely available [39]. In contrast, radar sensors, particularly millimeter wave (mmWave) radars [20], return only about a hundred points per frame and are noisier [15]. Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles [10].
Figure 1: Once an MDE model predicts incorrect relative depths between three or more objects, a linear scale-and-shift (dashed) cannot resolve this misalignment. Instead, we learn a polynomial fit with inflection points, enabling flexible non-uniform corrections.
@@
Unless camera baselines are known, images offer scaleless reconstruction. However, they are high-dimensional and sensitive to variations in illumination, object appearance, orientation, and camera viewpoint. Training robust models that can generalize across these factors often demands large-scale datasets, which are costly and time-intensive to collect. As a result, leveraging pretrained monocular depth estimation (MDE) models [41, 56, 4, 37] has emerged as a practical alternative. However, as these models only take in a single image, where reconstruction is inherently ill-posed, they typically infer scaleless relative depth, or metric depth that lacks the fidelity needed (see Tab. 2) for applications requiring accurate metric-scale reconstruction, e.g., navigation.
@@
Existing approaches [60, 51, 58, 18, 8, 59] attempt to transform these MDE predictions into accurate metric depth maps using a global linear (scale-and-shift) transformation. This class of methods assumes that the reconstruction is off by a single scaling factor across the entire scene. While effective in estimating ordinal relationships within local regions, where MDE predictions typically exhibit reasonable local (object-level) reconstructions, such linear corrections fail when multiple objects are placed at incorrect depths relative to each other. Specifically, once an MDE model places three or more objects at incorrect relative depths, no simple scale-and-shift can reconcile this misalignment (see Fig. 1). Empirically, a learned linear fit on a challenging dataset such as nuScenes performs worse than even a deterministic estimator like median scaling (see Tab. 2), indicating susceptibility to outliers.
@@
In this work, we challenge the common misconception that scale ambiguity in MDE is only up to an unknown global scale and shift. Instead, we propose to incorporate higher-order corrections through a polynomial transformation that allows for stretching and compressing at different depth levels, thereby enabling more flexible non-uniform adjustments. Our approach, PolyRad, transforms the predictions of pretrained MDE models by: learning prototypical patterns in the configuration of radar points, establishing spatial correspondences between radar and MDE features to encode a unified multi-sensor scene representation, and predicting polynomial coefficients that adaptively fit scaleless depth into metric depth. This polynomial expansion introduces additional degrees of freedom (i.e., multiple inflection points) that can better correct non-uniform variations and inter-region misalignments. While one may adjust every prediction (pixel) of an MDE model [25], the number of degrees of freedom is the total number of pixels. Given radar point clouds of a few hundred points (many orders less than the number of pixels), this becomes an ill-posed problem. On the other hand, the number of degrees of freedom in polynomial expansion is limited to the number of polynomial terms, which we empirically found to be close to the cardinality of radar point clouds, allowing our solution to be better posed and more regular. To our knowledge, no prior work has explored polynomial fitting for adapting predictions of pretrained foundation models, making our approach a novel solution not only for radar-guided depth estimation but also for any task that leverages foundation models and refines their predictions using additional inputs as guidance.
@@
However, learning such a flexible transformation is challenging because higher-order polynomials introduce a vast function space with many free parameters. Unlike a simple linear fit that relies on only scale and bias terms, polynomial fitting allows for numerous inflection points and nonlinearities. This expressiveness can inadvertently lead to harmful non-monotonic transformations—where the relative depth orderings of points are incorrectly reversed—if not properly constrained. We address this by introducing a regularization term that enforces the first derivative of the predicted metric depth with respect to the input MDE depth to generally remain close to 1, ensuring that incremental changes in the input MDE depth result in approximately equal changes in the predicted depth within local regions where MDE reconstructs structure with high fidelity. In essence, this regularizer encourages a smooth, monotonic transformation, mitigating unstable oscillations that can arise from overfitting high-degree polynomials while still enabling necessary corrections for cross-region misalignments.
@@
Our contributions: We propose (1) a novel radar-guided depth estimation method that leverages polynomial fitting to transform scaleless MDE predictions into accurate metric depth maps. Our approach (2) introduces a fundamental insight: using polynomial coefficients predicted from radar points to enable non-uniform, depth-dependent corrections. We present (3) a principled geometric approach, where polynomial transformations naturally introduce inflection points that can correct misalignments between local areas that a global scale-and-shift transformation cannot. We design (4) a novel training objective that encourages monotonicity through first-derivative as regularization, ensuring structural consistency while allowing inter-region depth adjustments. Finally, (5) extensive experiments on the nuScenes, ZJU-4DRadarCam, and View-of-Delft datasets demonstrate that PolyRad achieves the state of the art, outperforming existing methods by 11.3%, 54.6%, and 35.2%, respectively.
@@ 2.1 Monocular Depth Estimation
With recent advances in the scalability of neural networks [9, 5], monocular depth estimation (MDE) models can infer relative depth from a single image across diverse, unseen domains [22, 40, 55, 56, 37, 4], benefiting from large-scale datasets. Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation [16, 2], uncertainty estimation [38], optical flow [62], and visual odometry [46], to improve relative depth estimation. However, monocular metric depth estimation inherently suffers from scale ambiguity, as estimating absolute depth from a single image is an ill-posed problem. Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity [51]. One way to address this limitation is to incorporate range-sensing modalities such as lidar [21, 11], radar [45], or visual-inertial odometry [52].
2.2 Range-Camera Depth Estimation
@@
Metric-scale depth estimation often requires recovering a dense depth map from sparse range measurements, guided by complementary camera inputs. Image-guided depth completion, such as lidar-camera depth estimation, leverages the strengths of each modality: images offer dense visual context and structural priors, while sparse depth provides absolute metric scale to resolve depth ambiguity. Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers [32, 33], late fusion, which maintains independent branches for different sensors before merging outputs [57, 54, 50, 12, 43], and multi-scale fusion, which captures both local details and global scene structure [23]. U-Net-like architectures with CSPN-based refinement [6] have been widely used for coarse-to-fine depth completion [17, 27, 36], with further improvements from adaptive convolutions [7], deformable convolutions [36, 53], and transformer-based attention mechanisms [42, 61].
2.3 Radar-Camera Depth Estimation
@@
While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions [39]. In contrast, radar provides a cost-effective alternative [19], offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle [35, 47, 44]. The ubiquity of radar sensors [10] in existing automotive and robotic platforms further supports their integration into depth estimation pipelines. Leveraging radar for metric depth estimation not only reduces costs but also enhances the robustness and scalability of perception systems, making it an attractive choice for real-world deployment.
@@
Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity [45] of radar point clouds. [26] uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map. [29] leverages Doppler velocity and optical flow to associate radar points with image pixels. Other methods refine radar depth using neighborhood pixel associations [30], height-extension of radar points to address elevation ambiguity [28], or by extracting self-supervisory signals from the radar point cloud [13]. [45] introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion. [24] mitigates distribution artifacts using sparse supervision, while [48] employs a two-stage confidence-driven approach. [49] uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features. Like our method, [25] starts with an initial MDE prediction but refines it through quasi-dense scale estimation, followed by local scale adjustment, rather than through a learned scene-fitting approach.
@@
While these existing methods employ complex architectures to refine radar-derived spatial information, they all operate within the paradigm of completion or confidence-driven fusion. They often rely on multi-stage training, explicit radar-pixel association learning, or additional supervisory signals, greatly increasing model complexity. In contrast, PolyRad employs a streamlined yet powerful architecture that directly predicts polynomial coefficients from radar and MDE features, enabling accurate and efficient fitting of scaleless depth to metric scale. It bypasses the need for complex multi-stage systems that learn explicit correspondences while enabling flexible scene-adaptive depth corrections.
Figure 2: Overview of PolyRad. PolyRad transforms scaleless MDE predictions into metric depth using polynomial fitting guided by radar features. Learnable prototypes extract patterns in the configurations of radar point clouds and are used to compute aggregated radar features. The geometry-aware MDE features are fused with the radar features to yield a unified scene representation, which is used to predict coefficients for polynomial fitting, enabling non-uniform corrections that improve accuracy beyond simple scale-and-shift transformations.
@@ 3 Method Formulation
As MDE models are trained on millions of images, they are robust to typical visual nuisance variability, spanning from illumination changes and object appearances to viewpoint shifts. To circumvent the need to collect many large-scale synchronized radar-camera datasets, we begin with a frozen pretrained MDE model $M$, which serves as a learned geometric prior, to infer a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the image. We then encode the point cloud $C$ and the scaleless depth map $z$ separately. Through a fusion process that captures correspondences between them, we predict $N+1$ polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\cdots,\hat{c}_{N}\}$ that are used to transform $z$ into a metric depth map $\hat{d}$:
$\hat{d}\,=\,\hat{c}_{0}+\sum\nolimits_{i=1}^{N}\hat{c}_{i}z^{i}.$ (1)
@@ 3.1 Motivation
Existing methods [60, 51, 58, 18, 8, 59] refine MDE predictions by applying simple scale-and-shift transformations. However, this approach lacks the complexity needed to correct non-uniform variations from ground truth across different depths. A practical illustration of this shortcoming arises when multiple objects or local regions appear in the same scene: MDE may accurately infer the reconstruction within each object region, yet place them at incorrect relative depths with respect to each other. Such cross-object misalignments cannot be corrected by a single global scaling factor (see Fig. 1). This violates the assumption that the MDE reconstruction is up to an unknown global scale and shift.
@@
To address this, our polynomial fitting approach exponentiates the MDE prediction to higher powers and learns coefficients that transform it into a metric depth map. Lower-order coefficients (including scale and shift) capture the global scene layout, while higher-order coefficients focus on local depth adjustments, correcting cross-object misalignments and fine-grained MDE errors.
@@
An intuitive understanding of polynomial fitting can be drawn from the geometric perspective of depth transformations. A simple scale and shift operation, mathematically represented as $\hat{a}z+\hat{b}$, uniformly stretches or compresses the entire reconstruction by the same factor across all depth levels. While effective for coarse global corrections, such a linear transformation has zero inflection points, limiting its flexibility to address cross-region misalignments in relative depth predictions. In contrast, the total degree of our polynomial fitting method determines the maximum number of potential inflection points, where the curvature of the depth transformation can change. For a polynomial $f(z)=\sum_{i=0}^{N}\hat{c}_{i}\,z^{i}$, an inflection point $z^{*}$ occurs if
[…]

### v2

@@ Abstract
We propose POLAR, a novel radar-guided depth estimation method that introduces polynomial fitting to efficiently transform scaleless depth predictions from pretrained monocular depth estimation (MDE) models into metric depth maps. Unlike existing approaches that rely on complex architectures or expensive sensors, our method is grounded in a fundamental insight: although MDE models often infer reasonable local depth structure within each object or local region, they may misalign these regions relative to one another, making a linear scale and shift (affine) transformation insufficient given three or more of these regions. To address this limitation, we use polynomial coefficients predicted from cheap, ubiquitous radar data to adaptively adjust depth predictions non-uniformly across depth ranges. In this way, POLAR generalizes beyond affine transformations and is able to correct such misalignments by introducing inflection points. Importantly, our polynomial fitting framework preserves structural consistency through a novel training objective that enforces local monotonicity via first-derivative regularization. POLAR achieves state-of-the-art performance across three datasets, outperforming existing methods by an average of 24.9% in MAE and 33.2% in RMSE, while also achieving state-of-the-art efficiency in terms of latency and computational cost.
@@ 1 Introduction
Metric 3D reconstruction is critical for spatial tasks such as self-driving (Maier et al., 2012; Gupta et al., 2021), where it is necessary for one to perceive the structure of the 3D environment in order to navigate. In many such systems, multiple sensors—including cameras, lidar, and radar—provide complementary information about the 3D scene. While lidar sensors offer dense and precise point clouds, they are expensive and not widely available (Raj et al., 2020). In contrast, radar sensors, particularly millimeter wave (mmWave) radars (Iizuka et al., 2003), return only about a hundred points per frame and are noisier (Han et al., 2023). Yet, they offer key advantages: they are far more cost-effective and energy-efficient, robust to challenging environmental conditions, and ubiquitously equipped on modern vehicles (Eichelberger & McCartt, 2016).
@@
Unless camera baselines are known, images offer scaleless reconstruction. However, they are high-dimensional and sensitive to variations in illumination, object appearance, orientation, and camera viewpoint. Training robust models that can generalize across these factors often demands large-scale datasets, which are costly to collect. As a result, leveraging pretrained monocular depth estimation (MDE) foundation models (Ranftl et al., 2021; Yang et al., 2024; Bochkovskii et al., 2024; Piccinelli et al., 2025) emerges as a practical alternative. However, as monocular 3D reconstruction is inherently ill-posed, they typically infer scaleless relative depth, or depth that lacks the fidelity needed for applications demanding accurate metric-scale reconstruction, such as mapping and navigation.
@@
Existing approaches (Yin et al., 2023; Viola et al., 2024; Hu et al., 2024; Zeng et al., 2025; Ding et al., 2025; Yu et al., 2025) attempt to transform these MDE predictions into accurate metric depth maps using a global affine (scale-and-shift) transformation. This class of methods assumes that the reconstruction is off by a single scaling factor across the entire scene. While effective in estimating ordinal relationships within local regions, where MDE predictions typically exhibit reasonable local (object-level) reconstructions, such linear corrections fail when multiple objects are placed at incorrect depths relative to each other. Specifically, once an MDE model places three or more objects at incorrect relative depths, no simple scale-and-shift can reconcile this misalignment (Fig. 1).
Figure 1: If an MDE model predicts incorrect relative depths between three or more objects, an affine scale-and-shift (dashed) cannot resolve this misalignment. POLAR (solid) overcomes this limitation by learning an NN’th-order polynomial fit with up to N2N-2 inflection points.
@@
In this work, we challenge the common assumption that scale ambiguity in MDE is only up to an unknown global scale and shift. Instead, we propose to incorporate higher-order corrections through a polynomial transformation that allows for stretching and compressing at different depth levels, enabling more flexible non-uniform adjustments. Our approach, POLAR, transforms the predictions of pretrained MDE models by (i) learning prototypical patterns in the configuration of radar points, (ii) establishing spatial correspondences between radar and MDE features to encode a unified multi-sensor scene representation, and (iii) predicting polynomial coefficients that adaptively fit scaleless depth into metric depth. This polynomial expansion introduces additional degrees of freedom (i.e., multiple inflection points) that can better correct non-uniform variations and cross-region misalignments. While one may adjust every prediction (pixel) of an MDE model (Li et al., 2024a), the number of degrees of freedom is the total number of pixels. Given radar point clouds of a few hundred points (many orders less than the number of pixels), this becomes an ill-posed problem. On the other hand, the number of degrees of freedom in polynomial expansion is limited to the number of polynomial terms, which we empirically found to be close to the cardinality of radar point clouds, allowing our solution to be better posed and more regular. To our knowledge, no prior work has explored polynomial fitting for adapting predictions of pretrained foundation models.
@@
However, learning such a flexible transformation is challenging because higher-order polynomials introduce a vast function space with many free parameters. Unlike a simple linear fit that relies on only scale and bias terms, polynomial fitting allows for numerous inflection points and nonlinearities. This expressiveness can inadvertently lead to harmful non-monotonic transformations—where the relative depth orderings of points are incorrectly reversed—if not properly constrained. We address this by introducing a regularization term that enforces the predicted metric depth with respect to the input MDE depth to remain approximately monotonically increasing. This ensures that incremental changes in the input depth result in proportional changes in the predicted depth within local regions where MDE reconstructs structure up to a relative scale. In essence, this regularizer encourages a piecewise monotonic transformation, mitigating unstable oscillations that can arise from overfitting high-degree polynomials while still enabling necessary corrections for cross-region misalignments.
@@
Our contributions: We propose (1) POLAR, a novel POLynomial fitting method that leverages complementary radAR guidance to transform scaleless monocular depth into accurate metric depth. Our approach (2) introduces a fundamental insight: using polynomial coefficients predicted from a learned multimodal representation to enable non-uniform corrections. We present (3) a principled geometric formulation, where polynomial transformations introduce inflection points that can correct misalignments between local areas that an affine transformation cannot. We design (4) a novel training objective that encourages monotonicity through first-derivative as regularization, preserving local ordinality while allowing cross-region adjustments. Finally, (5) extensive experiments demonstrate that POLAR achieves the state of the art in both performance and efficiency, outperforming existing methods by an average of 29.1% and simultaneously delivering real-time processing of over 40 fps.
@@ 2 Related Work
Monocular Depth Estimation. With recent advances in the scalability of neural networks (Dosovitskiy et al., 2021; Caron et al., 2021), monocular depth estimation (MDE) models can infer scaleless relative depth from a single image across diverse, unseen domains (Ranftl et al., 2020; Yang et al., 2024; Piccinelli et al., 2025; Bochkovskii et al., 2024), benefiting from large-scale datasets. Recent works seek to enhance MDE by leveraging auxiliary image signals, such as structure and motion priors from segmentation (Hoyer et al., 2023; Bian et al., 2019), uncertainty estimation (Poggi et al., 2020), optical flow (Zhao et al., 2020), and visual odometry (Song et al., 2023), to improve relative depth estimation. However, monocular metric depth estimation inherently suffers from scale ambiguity, as estimating absolute depth from a single image is an ill-posed problem. Even MDE models trained with metric depth supervision often struggle to generalize to unseen domains with high fidelity (Viola et al., 2024). One way to address this limitation is to incorporate range-sensing modalities such as lidar (Jaritz et al., 2018; Ezhov et al., 2024), radar (Singh et al., 2023), or visual-inertial odometry (Wong et al., 2020).
@@
Image-Guided Depth Completion. Most often studied in the context of lidar-camera depth estimation, image-guided depth completion leverages the strengths of each modality: images offer dense visual context and structural priors, while lidar points provide absolute metric scale to resolve depth ambiguity. Fusion strategies for these modalities include early fusion, where feature maps are concatenated at initial layers (Ma & Karaman, 2018; Ma et al., 2018), late fusion, where inputs are processed by independent branches (Yan et al., 2021; Rim et al., 2025), and multi-scale fusion, which captures both local details and global scene structure (Li et al., 2020). U-Net-like architectures have been widely used for coarse-to-fine depth completion (Hu et al., 2021; Lin et al., 2022), with improvements from deformable convolutions (Park et al., 2020; Xu et al., 2020), and attention mechanisms (Rho et al., 2022; Zhang et al., 2023).
@@
Radar-Camera Depth Estimation. While lidar-based depth estimation methods achieve high accuracy due to their dense and precise measurements, their widespread adoption is limited by high costs, power consumption, and sensitivity to environmental conditions (Raj et al., 2020). In contrast, radar provides a cost-effective alternative (Hunt et al., 2024), offering robustness in adverse conditions such as low light, fog, and rain—where both cameras and lidar often struggle (Paek et al., 2022; Srivastav & Mandal, 2023). The ubiquity of radar sensors (Eichelberger & McCartt, 2016) in existing automotive and robotic platforms further supports their integration into depth estimation pipelines. Leveraging radar for metric depth estimation not only reduces costs but also enhances the robustness and scalability of perception systems, making it an attractive choice for real-world deployment.
@@
Despite these advantages, methods that fuse image and radar inputs for depth estimation must address the sparsity and elevation ambiguity (Singh et al., 2023) of radar point clouds. (Lin et al., 2020) uses a two-stage late fusion approach that first produces a coarse depth map and performs outlier rejection, then predicts the final depth map. (Long et al., 2021) leverages Doppler velocity and optical flow to associate radar points with image pixels. (Lo & Vandewalle, 2021) refines radar depth using height-extension of radar points to address elevation ambiguity. (Singh et al., 2023) introduces RadarNet, which uses radar-pixel correspondence scores as well as confidence scores to generate a semi-dense depth map, which is then used to predict the final dense depth map using gated fusion. (Li et al., 2024b) mitigates distribution artifacts using sparse supervision, while (Sun et al., 2024) employs a two-stage confidence-driven approach. GET-UP (Sun et al., 2025) uses attention-enhanced graph neural networks to capture both 2D and 3D features from radar data and leverages point cloud upsampling to refine radar features. Like our method, RadarCam-Depth (Li et al., 2024a) and TacoDepth (Wang et al., 2025) also begin with an initial MDE prediction, but refine it and decode dense depth directly from fused features, rather than employing a learned scene-fitting approach.
@@
Existing methods operate within the paradigm of completion or direct decoding of fused features, and they often rely on multi-stage training and explicit radar-pixel association learning, increasing model complexity and computational overhead. In contrast, POLAR employs a streamlined yet powerful architecture that directly predicts polynomial coefficients from radar and MDE features, enabling accurate and efficient fitting of scaleless depth to metric scale. It bypasses the need for multi-stage systems that learn explicit correspondences while enabling flexible scene-adaptive depth corrections.
@@ 3 Method Formulation
As MDE models are trained on millions of images, they are robust to typical visual nuisance variability, spanning from illumination changes and object appearances to viewpoint shifts. To circumvent the need to collect many large-scale synchronized radar-camera datasets, we begin with a frozen pretrained MDE model $M$, which serves as a learned geometric prior, to infer a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the image. We then encode the point cloud $C$ and the scaleless depth map $z$ separately. Through a fusion process that captures correspondences between them, we predict $N+1$ polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\cdots,\hat{c}_{N}\}$ that are used to transform $z$ into a metric depth map $\hat{d}$.
Figure 2: Method Overview. POLAR transforms scaleless MDE predictions into metric depth using polynomial fitting guided by radar features. Learnable prototypes extract patterns in the configurations of radar point clouds and are used to aggregate spatially-informed radar features. The geometry-aware MDE features are fused with the radar features via a learnable soft-correspondence module to yield a unified scene representation that is used to predict polynomial coefficients for fitting. This enables non-uniform corrections that improve accuracy beyond affine transformations.
@@ 3.1 Motivation
Existing methods (Zeng et al., 2025; Ding et al., 2025; Yu et al., 2025; Viola et al., 2024; Hu et al., 2024; Yin et al., 2023) refine MDE predictions by applying simple scale-and-shift transformations. However, this approach lacks the complexity needed to correct non-uniform variations from ground truth across different depths. A practical illustration of this shortcoming arises when multiple objects or local regions appear in the same scene: MDE may accurately infer the reconstruction within each object region, yet place them at incorrect relative depths with respect to each other. Such cross-region misalignments cannot be corrected by a single global scaling factor (see Fig. 1). This violates the assumption that the MDE reconstruction is up to an unknown global scale and shift.
@@
To address this, our polynomial fitting approach exponentiates the MDE prediction to higher powers and learns coefficients that transform it into a metric depth map via summation. Lower-order coefficients (including scale and shift) capture the global scene layout, while higher-order coefficients focus on local depth adjustments, correcting cross-object misalignments and fine-grained MDE errors.
@@
An intuitive understanding of polynomial fitting can be drawn from the geometric perspective of depth transformations. An affine (scale-and-shift) operation, mathematically represented as $\hat{a}z+\hat{b}$, uniformly stretches or compresses the entire reconstruction by the same factor across all depth levels. While effective for coarse global corrections, such an affine transformation has zero inflection points, limiting its flexibility to address cross-region misalignments in relative depth predictions. In contrast, the total degree of our polynomial fitting method determines the maximum number of potential inflection points, where the curvature of the depth transformation can change. For a polynomial $f(z)=\sum_{i=0}^{N}\hat{c}_{i}\,z^{i}$, an inflection point $z^{*}$ occurs if
$f^{\prime\prime}(z^{*})\,=\,\sum\nolimits_{i=2}^{N}i(i-1)\,\hat{c}_{i}\,(z^{*})^{i-2}\,=\,0,$ (1)